# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ Appendix H Proof of Proposition 7
For $|K|=3$, subtracting the third “block column” of the determinant from the first and second block columns, and then adding the first and second “block rows” to the third block row, the determinant in (H.1) becomes an upper triangular matrix:
$|\lambda I-J\Phi|=\begin{vmatrix}&\lambda I-(1-\alpha)I,&0&-X\\ &0&\lambda I-(1-\alpha)I&-Y\\ &0&0&\lambda I-X-Y-Z-(1-\alpha)I\\ \end{vmatrix}=0,$ (H.4)
@@
$\left|{\lambda I-J\Phi}\right|=\left|{\lambda I-(1-\alpha)I}\right|^{2}\cdot\left|{\lambda I-X-Y-Z-(1-\alpha)I}\right|=0.$ (H.5)
@@
$\left(\lambda_{i,1}-1+\alpha\right)\left(\lambda_{i,2}-1+\alpha\right)\left(\lambda_{i,3}-\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})\right)=0,$ (H.6)
@@
Referring to Theorem A.1, the system is stable if and only if all the $\lambda_{i}^{\prime}\text{s}\in(-1,1)$. Observing (H.2), (H.3) and (H.6), since $1-\alpha\in(0,1)$ when $10<\alpha<$ 1, the stability depends on $f(\rho_{i};\alpha)$, $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ and $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})$, respectively, for $|K|=1$, $|K|=2$ and $|K|=3$. Note that $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})$ contains $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ as a special case when $p^{2}=0$ (and thus $q_{2}^{1}=\frac{p^{1}}{p^{0}+p^{1}}=p^{0}$). Moreover, $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ contains $f(\rho_{i};\alpha)$ as a special case when $p^{0}=1$, $p^{1}=0$ and $\hat{\alpha}=\alpha$. Therefore, we can simply consult whether $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})\in(-1,1)$ or not to check the local stability.
Appendix I Proof of Proposition 8

### v2

@@ Appendix I Derivation of Fact 2
For $|K|=3$, subtracting the third “block column” of the determinant from the first and second block columns, and then adding the first and second “block rows” to the third block row, the determinant in Eq. (J.1) becomes an upper triangular matrix:
$|\lambda I-J\Phi|=\begin{vmatrix}&\lambda I-(1-\alpha)I,&0&-X\\ &0&\lambda I-(1-\alpha)I&-Y\\ &0&0&\lambda I-X-Y-Z-(1-\alpha)I\\ \end{vmatrix}=0,$ (J.4)
@@
$\left|{\lambda I-J\Phi}\right|=\left|{\lambda I-(1-\alpha)I}\right|^{2}\cdot\left|{\lambda I-X-Y-Z-(1-\alpha)I}\right|=0.$ (J.5)
@@
$\left(\lambda_{i,1}-1+\alpha\right)\left(\lambda_{i,2}-1+\alpha\right)\left(\lambda_{i,3}-\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})\right)=0,$ (J.6)
@@
Referring to Theorem A.1, the system is stable if and only if all the $\lambda_{i}^{\prime}\text{s}\in(-1,1)$. Observing Eqs. (J.2), (J.3) and (J.6), since $1-\alpha\in(0,1)$ when $10<\alpha<$ 1, the stability depends on $f(\rho_{i};\alpha)$, $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ and $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})$, respectively, for $|K|=1$, $|K|=2$ and $|K|=3$. Note that $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})$ contains $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ as a special case when $p^{2}=0$ (and thus $q_{2}^{1}=\frac{p^{1}}{p^{0}+p^{1}}=p^{0}$). Moreover, $g(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1})$ contains $f(\rho_{i};\alpha)$ as a special case when $p^{0}=1$, $p^{1}=0$ and $\hat{\alpha}=\alpha$. Therefore, we can simply consult whether $\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})\in(-1,1)$ or not to check the local stability.
Appendix K Proof of Proposition 9